OpenStudy (anonymous):

A rectangular prism has a volume of nine hundred cubic centimeters, and the base has an area of fifty square centimeters. What is the height of the prism?

OpenStudy (anonymous):

kk i can help

OpenStudy (anonymous):

b*h=900

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

we know b=50 so 50*h=900

OpenStudy (anonymous):

to find h we need to isolate it which means divide both sides with 50

OpenStudy (anonymous):

that makes h=18

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

also remember me?

OpenStudy (anonymous):

earlier in the morning ;3

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

you and the "squirt gun" XD

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

900 divided by 50 is 18 and that's your answer

OpenStudy (anonymous):

yeah we setteled that but thanks!

OpenStudy (anonymous):

can u help me with another one realy quick

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

The volume of a cone is one hundred twenty cubic meters. What is the volume of a cylinder with the same radius and the same height as the cone?

OpenStudy (anonymous):

k whats the formula to get its area?

OpenStudy (anonymous):

πr(r+h2+r2)

OpenStudy (anonymous):

h^2 and r^2 just saying.

OpenStudy (anonymous):

so 120=pi(r)(r+h^2+r^2)